# UniGenX: Unified Generation of Sequence and Structure with Autoregressive Diffusion

### v1

@@ Appendix D Computing Resource Usage
Table S2 details the GPU computational resource usage for various tasks, including task type, dataset, model size, GPU specifications, and computational cost (GPU days). All tasks were trained using float32 precision, as reflected in the table. Tasks were trained using either A100 (80G) or MI300X (192G) GPUs, with computational costs ranging from 0.9 GPU days (Carbon-24) to 191 GPU days (Unified Pretraining). While bf16 precision can achieve an approximate 4x training speedup without compromising accuracy, these results are not included in the table.
@@
Table S2: GPU Usage for Different Tasks
@@ E.5 Pretraining or Training from Scratch
Methods MR(%$\uparrow$ RMSD $\downarrow$ MR(%$\uparrow$ RMSD $\downarrow$ MR(%$\uparrow$ RMSD $\downarrow$

### v2

@@ Appendix D Computing Resource Usage
Table S2 details the GPU computational resource usage for various tasks, including task category, dataset, model size, GPU specifications, and total computational cost measured in GPU days. All experiments were conducted using either NVIDIA A100 (80G) or AMD MI300X (192G) GPUs. Unless otherwise specified, training was performed with BF16 precision for computational efficiency. The reported GPU days range from 0.9 (Carbon-24) to 191 (Unified Pretraining).
It is worth noting that during fine-tuning phases, smaller batch sizes were often used due to limited dataset sizes or task-specific constraints. As a result, GPU utilization was suboptimal, and the actual effective GPU compute required may be significantly lower than the reported GPU days if fewer GPUs were used in parallel.
@@
Conditional Mat Mag UniGenX(400M) A100 (80G) 55.8
Bulk UniGenX(400M) A100 (80G) 1.08
Multiple UniGenX(400M) MI300X (192G) 4.72
@@
Protein Pretrain UniGenX(400M) MI300X (192G) 853
MD UniGenX(400M) MI300X (192G) 11
EC Number UniGenX(400M) MI300X (192G) 41
Docking MISATO UniGenX(100M) MI300X (192G) 67.7
Table S2: GPU Training Cost per Task. Models trained with BF16 precision are underlined.
@@ E.5 Pretraining or Training from Scratch
[…]
@@
Appendix F Special Tokens Used in Sequence Modeling
In our model framework, we introduce a set of special tokens to indicate the beginning of different modalities. These tokens serve as explicit prompts for the model to distinguish between different input types, such as molecular SMILES, protein sequences, and geometric structures. Table S9 provides a summary of the special tokens used in our vocabulary.
Token Description
<bos> Beginning of sites/sentence/SMILES
<eom> End of sentence
<bof> Beginning of material formula
<bol> Beginning of material lattice
<bop> Beginning of protein amino-acid sequence
<boc> Beginning of coordinates
<bulk> Beginning of the property bulk modulus
<band> Beginning of the property band gap
<mag> Beginning of the property magnetic
<heat_capacity> Beginning of the property heat capacity
<E_hill> Beginning of the property energy above the convex hull
<density> Beginning of the property density
<bopo> Beginning of pocket
<eopo> End of pocket
<boapc> Beginning of pocket apo coords
<eoapc> End of pocket apo coords
<bom> Beginning of molecule SMILES sequence
<eom> End of molecule SMILES sequence
<bohpc> Beginning of pocket holo coords
<eohpc> End of pocket holo coords
<bomc> Beginning of molecule coords
<eomc> End of the molecule coords
<ec1> Beginning of the EC Number 1
<ec2> Beginning of the EC Number 2
<ec3> Beginning of the EC Number 3
Table S9: Explanation of special tokens used in the unified vocabulary.